5 peimts- To examine the relationship between the amount of questions on my statistics midterms and the time it takes to complete it, I gave 5 midterms # Questions Time (min 6 70 8 81 10 85 8 76 78 Hint: Σχ¡= 41, yi-390, У-345, Σ -30546, yiyi-3228 (a) State the regression equation. (p value can be within an interval) (b) At the 0.05 significance level, is the number of questions on an exam a significant predictor of the amount of time it takes to finish? Make sure to define the hypotheses. (c) IfI want an exam to last 90 minutes, how many questions should it be?
Answer #1

Let Y be the time (in mins) taken complete the midterm and X be the number of questions in the mid term.

The regression equation that we want to estimate is

Y=eta_0+eta_1X+epsilon

where eta_0 is the intercept of the regression line

61 is the slope of regression line

epsilon stackrel{iid}sim mathcal{N}(0,sigma^2) is a random error

We calculate the following

n=5

ŽZİ 41 8.2

390 78 =_== 78 390

SS_x=sum (x_i-ar{x})^2=sum x^2-nar{x}^2=345-5 imes 8.2^2=8.8

уг-ny2-30546-5x 782-126

SSr,- (zi-i) (Vi-у)- 2.iVi _ n.ry-3228-5x 8.2x 78-30

The estimate of slope is

S.S 30 3.409 8.8=3.409

The estimate of intercept is

30 = у-31-r = 78-3.409 × 8.2 = 50.045

a) The estimated regression equation is

y 50.045 3.409x

b) The number of questions on an exam (X) a significant predictor of the amount of time it takes to finish (Y), if the slope coefficient corresponding to X (61) is not equal to zero.

We want to test the following hypotheses

H0 : β1-0 ← null hypothesis: the number of questions is not a predictor of amount of time taken H : 31メ0 ← alternative hypoth

We need to calculate the following

Sum of square Errors is

egin{align*} SSE=SS_y-hat{eta}_1SS_xy=126-3.409 imes 30=23.727 end{align*}

Mean square Error is

egin{align*} MSE=rac{SSE}{n-2}=rac{23.727}{5-2}=7.909 end{align*}

The standard error of estimate is

egin{align*} s=sqrt{MSE}=sqrt{7.909}=2.8123 end{align*}

The standard error of the slope estimate is

egin{align*} s(hat{eta}_1)=rac{s}{sqrt{SS_x}}=rac{2.8123}{sqrt{8.8}}=0.9480 end{align*}

The hypothesized value of the slope is egin{align*} eta_{1H_0}=0 end{align*}

the test statistics is

egin{align*} t=rac{hat{eta}_1-eta_{1H_0}}{s(hat{eta}_1)}=rac{3.409-0}{0.9480}=3.596 end{align*}

The degrees of freedom is n-2=5-2=3

This is a 2 tailed test (The alternative hypothesis has "not equal to"). The p value of this statistics is

P(T<-3.596) +P(T>3.596)

Using the standard t tables, we can get the p-value with in an interval.

For 3 degrees of freedom, we can see that the total area under 2 tails of t-distribution is 0.05, when t=3.182 and is 0.02 when t=4.541.

The test statistics is 3.596. That means the p-value is in the interval 0.02 to 0.5

(We can get the exact p-value in excel using =T.DIST.2T(3.596,3) as 0.037)

We will reject the null hypothesis if the p-value is less than the significance level.

Here the p-value is in the interval 0.02 to 0.05, and hence is less than the significance level of 0.05.

Hence we will reject the null hypothesis.

We conclude that at 0.05 level of significance, the number of questions on an exam a significant predictor of the amount of time it takes to finish.

c) Using the equation

y 50.045 3.409x

we need to solve for X when the predicted value of Y is 90

90=50.045+3.409Ximplies X=rac{90-50.045}{3.409}=11.72

If I want an exam to last 90 minutes, there should be 12 questions in the exam.